# An Online Data Analysis Framework for Accelerator-Based Physics Experiments

### v2

@@ Abstract
A robust and flexible architecture capable of providing real-time analysis on diagnostic data is of crucial importance to physics experiments. In this paper, we present such an online framework, used in June 2025 as part of the HRMT-68 experiment, performed at the HiRadMat facility at CERN, using the Super Proton Synchrotron (SPS) beam line. HRMT-68 was a fixed-target laboratory astrophysics experiment aiming to identify plasma instabilities generated by a relativistic electron-positron beam during traversal of an argon plasma. This framework was essential for experimental data acquisition and analysis, and can be adapted for a broad range of experiments with a variety of experimental diagnostics. The framework’s modular and customizable design enabled us to rapidly observe and extract emergent features from a diverse range of diagnostic data. Simultaneously, it allowed for both the introduction of new diagnostic devices and the modification of our analysis as features of interest were identified. As a result, we were able to effectively diagnose equipment malfunction, and infer the beam’s response to varying bunch duration, beam intensity, and the plasma state without resorting to offline analysis, at which time adjustment or improvement would have been impossible. We present the features of this agile framework, whose codebase we have made publicly available, which can be adapted for future experiments with minimal modification.
@@
[nkua] organization=Department of Physics, National and Kapodistrian University of Athens, addressline=Zografos, city=Athens, postcode=GR-15771, country=Greece
@@
[lancaster] organization=Physics Department, Lancaster University, city=Lancaster, postcode=LA1 4YB, country=United Kingdom
@@ 1 Introduction and Motivation
Rapid data acquisition is imperative for modern experimental physics. With the advent of the High-Luminosity upgrade to the Large Hadron Collider (LHC) (HL-LHC) at CERN, it is estimated that ATLAS’s Trigger and Data Acquisition (TDAQ) systems will be capable of processing data at an input rate of [1]. Such extensive upgrades to data collection at CERN have already been introduced during the run-up to Run 3 of the LHC [2, 3], which began in 2022. Furthermore, experiments at CERN often rely on observing events of exceptional rarity, with the creation of the Higgs boson occurring only approximately once in every proton-proton collisions [4]. The data extracted from these events require extensive analysis, and the volume of data collection has often necessitated the use of sophisticated architectures, such as neural networks [5, 6].
@@
Even in smaller experiments, this scope of data collection, and the often-small fraction of shots occupied by events of interest, motivates the construction of software frameworks not only capable of discerning events of interest among a large quantity of data, but also of performing analysis and extracting features of interest during experimental runtime. This ability to perform data analysis in real-time is of immense value in cases where offline analysis would fail to identify faults such equipment failure or beam misalignment until long after the experiment’s conclusion.
@@
In this paper, we present such a robust, flexible, and reusable framework for online data analysis of experimental diagnostics, first employed during the HRMT-68 experiment performed at the HiRadMat facility [7] at CERN, part of the Super Proton Synchrotron (SPS) accelerator [8]. This was a laboratory astrophysics (see the discussion of recent advances in laboratory astrophysics by Savin et al. [9]) experiment, which aimed to study plasma instabilities [10, 11] produced by an electron-positron pair beam [12] traversing an argon plasma [13]. This framework places an emphasis on readability and modularity, making it easily-recyclable for a use in a diverse range of future experiments. Authored in Python, with the option to incorporate a JavaScript Object Notation (JSON) configuration file (see Section 4), it conforms rigidly to the object-oriented programming (OOP) paradigm for maximized encapsulation and abstraction [14, 15], producing a software package that can be easily utilized by those with minimal programming expertise. The resulting framework executes most of its complex data extraction and analysis pipeline without directly interfacing with the user; this “under-the-hood” operation is the framework’s primary strength.
@@
The paper is outlined as follows: the motivation and experimental layout of the HRMT-68 experiment is discussed in Section 2. We then present our methods of continuous data acquisition in Section 3, which was crucial in developing a central repository of raw data for online analysis during the experimental run. In Section 4, we discuss the code itself, outlining its hierarchical structure, modular design, and robustness, before quantifying its performance in Section 5. The paper concludes with a discussion of the framework’s strengths and weaknesses, as well as any potential improvements for future incarnations of the code, in Section 6.
@@ 2 The HRMT-68 Experiment
The HRMT-68 experiment utilizes a novel platform recently developed at CERN’s HiRadMat facility, which generates an electron-positron pair-beam from the collision of a -momentum proton beam with a fixed-target, as detailed by Arrowsmith et al. [12, 13, 16]. This electron-positron pair-beam passes into a cylindrical argon plasma, produced via inductive discharge [17, 16, 18].
@@
The complexity of this experiment required a diverse collection of diagnostic equipment, which, collectively, had to provide optical, spectroscopic, and magnetic field data from the plasma. A summary of the devices used, and their purpose, is provided in the form of Table 1, with specific discussion of the individual devices’ purpose and function(s) provided in Sections 2.1-2.4.
2.1 B-dot Probe Oscilloscopes
Magnetic field diagnostics for the plasma were provided by three B-dot probes, named for the overdot denoting differentiation of the magnetic field with respect to time: [19]. These probes measure the magnetic field strength via a practical application of Faraday’s Law [20],
where is the e.m.f. induced in the coil, is the number of turns in the B-dot probe coil, is the observed magnetic field vector, and represents an infinitesimal vector area element of a single turn of the coil. Magnetic field strength can thus be inferred via time-integration of the induced voltage in the coil (for a more detailed derivation, we point the reader to the work of Bose et al. [19]).
The e.m.f. induced in the B-dot coils was measured using two four-channel oscilloscopes (Tektronix 6 Series MSO [21]). Of the eight oscilloscope channels, seven were used to collect data from the B-dot probes, with the eighth channel left unused.
Two oscilloscope channels were allocated to each B-dot probe in order to distinguish between inductive and capacitive signals; a relative coil rotation with respect to the magnetic field lines is introduced between the two channels, resulting in induced voltages of opposite sign. Therefore, subtraction between each probe’s two channels is expected to produce a nonzero signal for inductive voltages, and zero for capacitive voltages, and formed a key part of the analysis.
Essential also for analysis was the extraction of both time and voltage amplitude data (“samples”) from the oscilloscope outputs. Time data included the sample number () and time step () between samples, in addition to the timestamp of each sample. This was used to calculate the Fast Fourier Transform (FFT) of each channel’s voltage data in order to image the magnetic field in both the time and frequency domains.
2.2 Chromox and Radiation-Hard Cameras
Visual diagnostics were provided by four screens coupled with optical cameras. Additionally, one extra radiation-hard camera provided a live feed of the plasma. This device produced no quantitative data, but was instead used to identify any plasma instabilities during plasma generation and therefore was not included in the online data analysis framework. Due to its role as a largely heuristic diagnostic, the radiation-hard camera’s images were not collected or saved, and no analysis was performed on this device’s output.
Four screens made of chromox, a chromium-doped alumina ceramic ( Al2O3, Cr2O3) [16, 22, 23], were used in the setup. Chromox produces scintillation light in the visible spectrum at wavelengths of and [22]. Of the four chromox screens used in the setup, two were associated with the cell, with one placed in the cell’s most downstream crosspiece, and the other in the cell’s most upstream crosspiece. The pair of cameras imaging these two screens are henceforth referred to as the “plasma cell cameras”. The data from the cameras were collected by the online framework, and first (Equation (2)) and second moment (Equation (3)) calculations were carried out online via weighting each pixel’s coordinates by its intensity, and dividing by the total summed intensity across all pixels, using Equation (2):
where and are the horizontal and vertical components, respectively, of the pixel’s coordinate vector denoting displacement from the origin (here taken to be the center of the image). These values used to quantitatively characterize the beam’s transverse cross-section.
Plots of intensity as a function of the Cartesian and coordinates were produced by summing over and , respectively.
where is the total number of pixels, is intensity of the pixel, and is the pixel coordinate. In Equation (3), denotes the element-wise product between two matrices.
Two additional sums were performed over the polar radial (), and azimuthal () coordinates of the image to produce plots of intensity as a function , and , respectively. values were calculated from the norm of each pixel’s coordinate vector, i.e. and where . and coordinates were quantized and placed in a finite number of and bins to produce smooth histogram plots of intensity versus both radial and azimuthal coordinates.
Two further chromox screens were introduced outside of the cell, beyond the dipole magnet: these screens were used to infer the electron and positron energy spectra of the beam after exiting the plasma cell. Thus, these two cameras were referred to as the “spectrometer cameras”. For these cameras, the code was designed to perform background-subtraction, ensemble averaging, and integrals over a specified axis.
2.3 Streak Unit Camera
@@
For any initial velocity, charged particles emit optical transition radiation (OTR) when passing between two media of differing dielectric constant; the variation in electric and magnetic fields across this boundary results in this radiative emission [24, 25]. The collection of photons produced via OTR is often used in beam diagnostics to characterize the longitudinal density profile of accelerator beams [26, 27]. OTR was collected by a digital camera mounted to a Hamamatsu streak unit [28] with an adjustable slit width, producing data which can be encoded in a 2D image containing information about the pulse duration of incident radiation. In addition to background-subtraction and ensemble averaging, the code was designed to allow integrals over a single axis of the image. These were produced by summing across the axis of the image perpendicular to the temporal axis, as per Equation (4).
where is the number of pixels corresponding to the image’s -dimension (column number), and is a matrix of dimensions image pixels, with each element of the matrix corresponding to pixel intensity. This produces , a matrix, with denoting the image’s -dimension and hence the number of time steps captured in the streak image, and where each entry in is therefore an approximate radiation intensity at each time step. is then used to produce an intensity versus time plot. The code was therefore constructed to both produce these plots, and a visual plot of the 2D image captured by the camera attached to the streak unit.
2.4 Synchrotron Spectrometer Camera
@@
Synchrotron radiation is produced by the deflection of charged particles in magnetic fields [29], and can be used to infer the interaction of the electron-positron beam with the magnetic fields produced in the plasma, and is imaged by a Andor iStar intensified charge-coupled device (CCD) camera [30] mounted to a mounted to a Czerny-Turner monochromator [31, 32].
The images produced by this camera are 2-dimensional, with one axis corresponding to wavelength of the collected radiation; this is due to the dispersion in the incident light produced by the spectrometer’s grating. The second axis encodes intensity of incident radiation at each wavelength at each pixel in the active region of the CCD. A Fourier transform can thus be produced by summing along all pixel elements for a given wavelength, i.e. along the dimension perpendicular to the axis encoding wavelength information. The code was designed to perform this function with or without background correction. Once more, the code was also refactored to allow statistical averaging over several images from the camera, and display statistical variations111Typically . in the Fourier Transforms produced by summing over one of the image’s dimensions.
@@
Device Purpose Chromox Cameras Plasma Cell Cameras Observe transverse profile of electron-positron pair-beam and/or proton beam. Spectrometer Cameras Observe energy spectrum of electrons + positrons exiting plasma cell. OTR Streak Unit + Camera Produce spatially- and temporally-resolved information about OTR produced by pair-beam. Synchrotron Emission Spectrometer + Camera Resolve wavelengths of radiation present in synchrotron emission due to electron-positron beam’s deflection by plasma B-field. B-dot Probes + Oscilloscopes Observe/plot B-field induced in plasma.
@@ 3.1 Construction of a Centralized Data Repository
CERN’s sophisticated Accelerator Data Logging Service (NXCALS) [33], which necessitates the existence of a special front-end, designated “Front-End Software Architecture” (FESA)222See https://confluence.cern.ch/spaces/viewspace.action?key=FESA3., was developed for LHC and its injectors, and is extremely versatile and adaptable. However, the versatile hardware used by this service, designed to accommodate a wide berth of accelerator-based experiments, was not sufficiently specialized for implementation in HRMT-68. This necessitated a system of logging and analysis external to the FESA/NXCALS framework.
This motivated the construction of an easily-accessible, experiment-specific central repository of data collected during the HRMT-68 campaign. In our case, we profited from CERN’s online Jupyter environment, Service for Web-based ANalysis (SWAN, described in the CERN SWAN documentation [34]), which allows users to run code from any location, with direct access to CERN’s EOS Open Storage (EOS)333EOS acronym is recursive.-CERNbox online file storage configuration; EOS is CERN’s disk storage system, with a storage capacity of [35], and CERNbox is CERN’s cloud-based file-sharing and storage service [36]. Data from each diagnostic was assigned a unique directory in this central data repository, which was pointed to by the code’s configuration (as described in Section 4). However, this framework would work equally well with any cloud-based or local file storage service.
In order to sync device data to the experiment’s CERNbox repository, each device was configured to either save data directly into CERNbox via a mounted network drive, or to a local directory synced continuously with the appropriate device directory on CERNbox, depending on device. The speed of the synchronization process varied greatly between devices, from seconds for devices with access to a mounted network drive, to seconds for devices whose local directories were synced remotely.

### v3

# An Online Data Analysis Framework for Small-Scale Physics Experiments

@@ Abstract
A robust and flexible architecture capable of providing real-time analysis on diagnostic data is of crucial importance to physics experiments. In this paper, we present such an online framework, used in June 2025 as part of the HRMT-68 experiment, performed at the HiRadMat facility at CERN, using the Super Proton Synchrotron (SPS) beam line. HRMT-68 was a fixed-target laboratory astrophysics experiment aiming to identify plasma instabilities generated by a relativistic electron-positron beam during traversal of an argon plasma. This framework was essential for experimental data acquisition and analysis, and can be adapted for a broad range of similar-scale experiments with a variety of experimental diagnostics, even those without a standard direct network communication interface. The developed framework’s customizable design enabled us to rapidly observe and extract emergent features from a diverse range of diagnostic data. Simultaneously, its modularity allowed for a quick introduction of new diagnostic devices and the modification of our analysis as features of interest were identified. As a result, we were able to effectively diagnose equipment malfunction, and infer the beam’s response to varying bunch duration, beam intensity, and the plasma state without resorting to offline analysis, at which time adjustment or improvement would have been impossible. We present the features of this agile framework, whose codebase we have made publicly available so that it may be adapted for future experiments with minimal modification.
@@
[liverpool] organization=Department of Physics, University of Liverpool, addressline=Brownlow Hill, city=Liverpool, postcode=L69 7ZX, country=United Kingdom
@@
[lancaster] organization=Department of Physics, University of Lancaster, city=Lancaster, postcode=LA1 4YB, country=United Kingdom
@@ 1 Introduction and Motivation
Small-scale accelerator experiments often face constraints from limited beam time and resources. As a typical example, the experiments in the CERN’s North Experimental area are allocated 1 week of total beam time, which includes setup, installation, beam tuning, data taking, and decommissioning [Banerjee:2774716]. Furthermore, experiments not undertaken at state-of-the-art facilities, for example the European Organisation of Nuclear Research (CERN), will not have access to specialist frameworks, such as CERN’s Accelerator Data Logging Service (NXCALS) [sobieszek23:1494]. Smaller experimental facilities which typically accommodate test-beam users and experiments, among them INFN-LNF [Balerna2023] or NCSR Demokritos [NCSRDemokritos_overview], lack the NXCALS infrastructure, increasing reliance on custom and dedicated equipment to perform online and offline data analysis. Even when such frameworks are present, they may be insufficiently flexible to address experimental needs. In novel platforms like the newly established FIREBALL platform at CERN [arrowsmith21:023103], experimental demands often change drastically during runtime, and frameworks that are flexible and capable of addressing these changes are essential.
Though there are conceptually more sophisticated and more reliable frameworks in operation, such as CERN’s JCOP/WinCC Open Architecture [GEMDCS_JCOP_Overview], we have chosen to instead deploy our framework during experimental runtime in order to maintain unlimited control over its functionality, code, and integration. This was essential for continuity between incarnations of the FIREBALL experimental platform, providing us with a framework tailored to our particular experimental requirements, and allowing internal continuity and collaborative modification between subsequent iterations of the experiment. Additionally, our framework was relatively quick to construct, minimally complicated, and highly customizable, and we modified it with ease even during our allocated beamtime to address issues ranging from a camera’s optical misalignment to equipment malfunction. Such malfunctions were common and often critical; our Czerny-Turner monochromator [czerny30:792], tasked with detecting synchrotron emission from within the plasma, ceased operation during experimental runtime. Though the device registered as still operational, it failed to capture images when triggered, and the cause of this failure remains unknown. Additionally, due to damage to an optic fibre, we could not deploy one of our optical components as planned. The code’s modular structure, wherein devices can be trivially added or removed from the data extraction protocol, allowed the framework to continue running unimpeded despite these types of failure. On our small collaboration, such obstacles are difficult to resolve due to the lack of technical expertise with respect to specialist equipment.
HRMT-68 differed from its sister experiments at CERN in that it collected a comparatively small amount of data, approximately in total. This is in sharp contrast to LHC-centred experiments. For example, during the LHC Run 3, LHCb collected raw data at a rate of [Fontana2025]. Similarly, ATLAS’ Trigger and Data Acquisition (TDAQ) systems will be capable of processing data at an input rate of after the High Luminosity upgrades at the LHC (HL-LHC) [Kopeliansky2023]. Our far smaller scope of data acquisition affected our design considerations, motivating the use of Python, whose performance in our framework is modest and would not be scalable to larger collaborations’ experimental needs.
@@
Particularly in small- or medium-size experiments, the scope of data collection, and the often relatively small fraction of beam spills (or pulses) corresponding to events of interest, motivates the development of software frameworks not only capable of discerning events of interest among a moderate quantity of data, but also of performing analysis and extracting features of interest during experimental runtime. This ability to perform data analysis in real-time is of immense value in cases where offline analysis would fail to identify faults such equipment failure or beam misalignment until long after the experiment’s conclusion.
@@
In this technical note, we present such a robust, flexible, and reusable framework for online data analysis of experimental diagnostics, first employed during the HRMT-68 experiment performed at the HiRadMat facility [efthymiopoulos11:1665] at CERN, part of the Super Proton Synchrotron (SPS) accelerator [arnaldi2025futurefacilitiescernsps]. This was a laboratory astrophysics (see the discussion of recent advances in laboratory astrophysics by Savin et al. [savin12:036901]) experiment, which aimed to study plasma instabilities [PhysRevLett.31.1390, Schekochihin2024] produced by an electron-positron pair beam [arrowsmith24:5029] traversing an argon plasma [arrowsmith21:023103]. Our developed framework places an emphasis on readability and modularity, making it easily applicable and reusable for a diverse range of future experiments. Authored in Python, with the option to incorporate a JavaScript Object Notation (JSON) configuration file (see Section 4), it conforms rigidly to the object-oriented programming (OOP) paradigm for maximized encapsulation and abstraction [lutz11b, phillips2015python], producing a software package that can be easily utilized by those with minimal programming expertise. The resulting framework executes most of its complex data extraction and analysis pipeline without directly interfacing with the user; this “under-the-hood” operation is one of the framework’s primary strengths.
@@
The present work is outlined as follows: the motivation and experimental layout of the HRMT-68 experiment is discussed in Section 2, provided to the reader as an exemplar application. We then present our methods of continuous data acquisition in Section 3, which was crucial in developing a central repository of raw data for online analysis during the experimental run. In Section 4, we discuss the code itself, outlining its hierarchical structure, modular design, and robustness, before quantifying its performance in Section 5. The paper concludes with a discussion of the framework’s strengths and weaknesses, as well as any potential improvements for future incarnations of the code, in Section 6.
@@ 2 The HRMT-68 Experiment
The HRMT-68 (FIREBALL-III) experiment utilizes a novel platform recently developed at CERN’s HiRadMat facility, which generates an electron-positron pair-beam from the collision of a -momentum proton beam with a fixed-target, as detailed by Arrowsmith et al. [arrowsmith24:5029, arrowsmith21:023103, Arrowsmith2024thesis]. This produced electron-positron pair-beam passes into a cylindrical argon plasma, produced via inductive discharge [arrowsmith23:P04008, Arrowsmith2024thesis, lieberman25b]. This plasma was contained in a glass plasma cell.
@@
The complexity of this experiment required a diverse collection of diagnostic equipment, which, collectively, had to provide optical, spectroscopic, and magnetic field data from the plasma. A summary of the devices used, and their purpose, is provided in the form of Table 1. For details on the setup on previous incarnations of the experiment, we point the reader again to refs. [arrowsmith24:5029, arrowsmith21:023103, Arrowsmith2024thesis].
Four screens made of Chromox, a chromium-doped alumina ceramic ( Al2O3, Cr2O3) [Arrowsmith2024thesis, Burger2016, McCarthy2003], were used in the setup. Chromox produces scintillation light in the visible spectrum at wavelengths of and [Arrowsmith2024thesis]. Each of the four screens was imaged individually using four cameras. These four cameras, designated , were collectively dubbed the “Chromox cameras”. Cameras imaged the plasma cell directly, and were referred to as the “plasma cell” cameras; the two other Chromox cameras imaged screens on the far side of a dipole magnet, and were dubbed the “spectrometer cameras”. Deflection of particles in a dipole field is a function of their energy, and hence scintillation patterns produced by deflected electrons/positrons as they impinge these two latter screens provide a measure of the electron/positron energy spectrum. Two screens were necessary, as electrons and positrons are deflected in opposite directions by a dipole magnetic field.
@@
For any initial velocity, charged particles emit optical transition radiation (OTR) when passing between two media of differing dielectric constant; the variation in electric and magnetic fields across this boundary results in this radiative emission [otr_dasilva, otr_gitter]. The collection of photons produced via OTR is often used in beam diagnostics to characterize the longitudinal density profile of accelerator beams [Chen2023, Fedorov_2020], as was done in the HRMT-68 run. OTR was produced by introducing a copper foil into the path of the beam, with the subsequent emitted radiation collected by a “streak unit” [hamamatsu] with an adjustable slit width. The information can be encoded in a 2D image, which contains information about the pulse duration of incident radiation, and is captured by a digital camera; such a digital camera was mounted on the streak unit during the experiment.
@@
Synchrotron radiation is produced by the deflection of charged particles in magnetic fields [Walker1994]. In the HRMT-68 experiment, synchrotron emission was used to infer the interaction of the electron-positron beam with the magnetic fields produced in the plasma. Synchrotron radiation emission from the plasma was collected using a charge coupled device (CCD) [Hui2020] camera (Andor iStar [andor_istar]) mounted on a Czerny-Turner monochromator [czerny30:792].
Magnetic field diagnostics for the plasma were provided by three B-dot probes, named for the overdot denoting differentiation of the magnetic field with respect to time: [Bose_2018]. These probes measure the magnetic field strength within the plasma via a practical application of Faraday’s Law: electromotive force (e.m.f.) induced in the coils of the B-dot probe is proportional to the rate of change of magnetic flux through the coils, using Equation 1 (see refs. [Griffiths_2017]).
The e.m.f. induced in the B-dot coils was measured using two identical four-channel oscilloscopes (Tektronix 6 Series MSO [textronix6]).
@@
Device Purpose Chromox Cameras Plasma Cell Cameras Observe transverse profile of electron-positron pair-beam. Spectrometer Cameras Observe energy spectrum of electrons & positrons exiting plasma cell. OTR Streak Unit + Camera Produce spatially- and temporally-resolved information about OTR produced by pair-beam. Synchrotron Emission Spectrometer + Camera Resolve wavelengths of radiation present in synchrotron emission due to electron-positron beam’s deflection by plasma B-field. B-dot Probes + Oscilloscopes Observe/plot B-field induced in plasma.
@@ 3.1 Construction of a Centralized Data Repository
CERN’s sophisticated Accelerator Data Logging Service (NXCALS) [sobieszek23:1494] necessitates the existence of a special front-end, designated “Front-End Software Architecture” (FESA)111See https://confluence.cern.ch/spaces/viewspace.action?key=FESA3.. FESA was developed for LHC and its injectors, and is extremely versatile and adaptable. However, its implementation requires specialist expertise usually not present on small-scale international collaborations, and many experiment-specific pieces of equipment are not trivially compatible with FESA. For example, the HRMT-68 used a NAVIO matching network [AdvancedEnergy2018Navio] for impedance matching when delivering RF power to the plasma cell; integrating the matching network into FESA would have required time and expertise that was unavailable to the collaboration.
Finally, it is generally known that the development of these generalized software frameworks (e.g. NXCALS/FESA) is driven by the global needs of each organization and its flagship experiments, leaving little to no space (in terms of resources or support) for the requirements and the versatility necessary for smaller experiments to be taken into account. This was our primary motivation for this work.
The variety of data from a range of diagnostics (Table 1) motivated the construction of an easily-accessible, experiment-specific central repository of data collected during the HRMT-68 campaign. Similar small-scale experiments may only have access to a generic cloud-based storage service accessible over a network connection, as was the case on HRMT-68. Therefore, our framework was designed to cater to this limitation.
In our case, we profited from CERN’s online Jupyter environment, Service for Web-based ANalysis (SWAN, described in the CERN SWAN documentation [swan_docs]), which allows users to run code from any location. The Jupyter environment provides a user-friendly front-end for executing Python without the need for SSH connections to remote servers or installation of local versions of Python. SWAN is directly integrated with CERN’s EOS Open Storage (EOS)222EOS acronym is recursive./CERNbox online file storage configuration; EOS is CERN’s disk storage system, with a storage capacity of [EOS_docs], and “CERNbox” is CERN’s cloud-based file-sharing and storage service [CERNbox_docs]. Therefore, the use of SWAN permitted any user with EOS/CERNbox read and write permissions and an internet connection to run our software from any location.
Data from each diagnostic was assigned a unique directory in this central data repository, which was pointed to by the code’s configuration (as described in Section 4). However, this framework would work equally well with any cloud-based or local file storage service.
In order to synchronize the different device data to the repository, each device was configured to either save data directly into CERNbox via a mounted network drive, or to a local directory synced continuously with the appropriate device directory on CERNbox, depending on the device. The speed of the synchronization process varied greatly between devices, from seconds for devices with access to a mounted network drive, to seconds for devices whose local directories were synced remotely.